OpenStudy (anonymous):

Integral using parts of t^(2) sin(t) I got it to -t^2cos(t) dt+2Integral tcos(t) dt How and why do I have to use parts again to solve this? Thanks

OpenStudy (kinggeorge):

So you've gotten to the point \[\int t^2\sin(t) dt=-t^2\cos(t)+2\int t\cos(t) dt\]To finish solving this, you need to integrate \[\int t\cos(t) dt\] using parts in the same way that you did the first integral.

OpenStudy (kinggeorge):

You have to use parts to solve it, because it's the best way of solving it after guessing.

OpenStudy (anonymous):

Why only integrate that side?

OpenStudy (kinggeorge):

What exactly do you mean?